OpenStudy (anonymous):

\[Calculate: \int\limits_{0}^{1/2}4x/\sqrt{1-x^2}dx\]

OpenStudy (anonymous):

I'm aware that I should substitute. The answer I come up with is - according to the book - not correct.

OpenStudy (anonymous):

\[-4+4\sqrt{3/4}\] Is what I came up with.

OpenStudy (anonymous):

try\[x ^{2} = t\] and dt =2xdx

OpenStudy (anonymous):

Trying!

OpenStudy (anonymous):

\[2\int\limits_{0}^{1/4}dt/\sqrt{1-t}\]

OpenStudy (anonymous):

I got -4+2*sqrt(3/4) Which according to the book is wrong, did I go wrong somewhere?

OpenStudy (accessdenied):

what does the book say is correct?

OpenStudy (anonymous):

4-2sqrt(3)

OpenStudy (experimentx):

if anyone has doubt about answer you can always check at wolframalpha

OpenStudy (anonymous):

I can't, don't know how to write the integrals! :c

OpenStudy (anonymous):

i think my answer should be correct

OpenStudy (experimentx):

let's ask wolf abt that .. hehe

OpenStudy (accessdenied):

"integral from 0 to 1/2 of 4x/sqrt(1-x^2) dx" wolfram agrees with book

OpenStudy (anonymous):

Blargh! :x

OpenStudy (experimentx):

accessdenied should have post the link instead

OpenStudy (anonymous):

i would try substituting u = 1 - x^2

OpenStudy (anonymous):

That's what I originally did eigenschmeigen, did not work out. I think I got the same result I do now, not sure.

OpenStudy (accessdenied):

true http://www.wolframalpha.com/input/?i=integral+from+0+to+1%2F2+of+4x%2Fsqrt%281-x%5E2%29+dx i dont really use this much, forgot about that. :P

OpenStudy (experimentx):

@eigenschmeigen you always point out that ... now i will never forget that.

OpenStudy (anonymous):

integral that i wrote befor value is: \[2[-2\sqrt{(1-t)}]_{0}^{1/4}=2(2-2\sqrt{3/4}) =4 -2\sqrt{3}\]

OpenStudy (anonymous):

How did the -2 get before the sqrt?

OpenStudy (anonymous):

That's where I'm failing..

OpenStudy (anonymous):

\[\int\limits (1-t)^{-1/2}\]

OpenStudy (anonymous):

and -t ofcource

OpenStudy (anonymous):

Ah, brainfart, thanks alot!

OpenStudy (anonymous):

ya, it's ugly

OpenStudy (accessdenied):

Thanks, myko! :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

\[\int\limits^{1/2}_0{\frac{4x}{\sqrt{1-x^2}}dx}\] \[u= 1-x^2\] \[\frac{du}{dx}= -2x\] \[\frac{-du}{2x}= dx\] \[\int\limits^{1/2}_0{\frac{4x}{\sqrt{1-x^2}}dx}=\int\limits^{3/4}_1{\frac{-2}{\sqrt{u}}du}= -4[\sqrt{u}]^{3/4}_1 = 4 - 2\sqrt{3}\] sorry if the question is already resolved - once i started i didnt want to stop

OpenStudy (anonymous):

Don't be sorry - thanks for helping! I know what you mean by not wanting to stop, absolutely loving this chapter regarding substitution. Mathematically intoxicating. :3

OpenStudy (anonymous):

absolutely agree, gotta love substitution